# Critical spin models from holographic disorder

### v1

@@ Acknowledgments
We wish to thank Lennart Bittel and Tommaso Guaita for fruitful discussions on Gaussian states, Jens Eisert for helpful comments on boundary conditions in critical spin chains, and Zoltán Zimborás for relevant discussions on aperiodic disorder. For all tensor network numerical simulations done in this paper, we used the QSpace tensor library Weichselbaum (2012, 2020, 2023) which can exploit general non-Abelian symmetries in tensor network algorithms. The authors are grateful for support from the Einstein Research Unit “Perspectives of a quantum digital transformation”.
@@ Appendix A Covariance matrix formalism and entanglement entropy calculation
A fermionic Gaussian state Surace and Tagliacozzo (2022) can be defined as any state
@@
The above expression reveals that the fermionic Gaussian state $\rho$ is completely characterized by the occupation numbers $\langle\tilde{c}^{\dagger}_{k}\tilde{c}_{k}\rangle$, and as a consequence by the correlators $\langle c_{i}^{\dagger}c_{j}\rangle$ and $\langle c_{i}c_{j}\rangle$ after we apply the reverse transformation, $c_{k}=U^{\dagger}\tilde{c}_{k}$. It is convenient to study a Gaussian model in the Majorana fermion picture. In this case the Gaussian fermionic state $\rho$ is completely characterized by the covariance matrix
@@
where the density matrix is a thermal density matrix and $\gamma_{k}$ are the Majorana operators with $\{\gamma_{j},\gamma_{k}\}=2\delta_{j,k}$. The covariance matrix $\Gamma$ is a real, skew-symmetric matrix, which makes its diagonalization practical: For any real, skew-symmetric $2N\times 2N$ matrix $M$, there exist a real special orthogonal matrix $O\in SO(2N)$, such that $M=O\tilde{M}O^{\text{T}}$, where $\tilde{M}$ is a block diagonal matrix of the form
@@
where $\mu_{j}\in\big{[}-\frac{1}{2},\frac{1}{2}\big{]}$. The eigenvalues $\epsilon_{k}$ of the Hamiltonian $H$ are then given by:
@@
Having obtained the eigenvalues $\mu_{j}$, the von Neumann entanglement entropy is given by:
$S(\rho)=\sum_{k=1}^{N}\big{[}\nu_{k}\ln(\nu_{k})+(1-\nu_{k})\ln(1-\nu_{k})\big{]}$ (17)
where $\nu_{k}=\frac{1}{2}-\mu_{k}$. The block diagonal form of the covariance matrix makes the calculation of the entanglement entropy of subsystems of any length $\ell$ quite convenient.
Figure 8: Site-averaged entanglement entropy $S_{\text{avg}}$ for subsystems of size $\ell=29$ over disorder strength for (a): hyperbolic sequences and (b): non-hyperbolic aperiodic sequences. In both (a) and (b) we present the result of the $\{3,7\}$-hyperbolic lattice with a black line for comparison. (c): Averaged entanglement entropy for subsystems of size $\ell=29$ over disorder strength for the Fibonacci and Silver-mean aperiodic sequences. (d): We present the behavior of the averaged entanglement over disorder strength for the 6th inflation step of the $\{3,7\}$-hyperbolic sequence and the 7th inflation step of the Silver-mean sequences. This is asymptotically the behavior of the averaged entanglement entropy for large number of inflation steps for these two cases.
@@ Appendix B Benchmarking DMRG
We can check the accuracy of the DMRG method we used for the disordered non-Gaussian models in our analysis above, by applying it also to the Gaussian model. For the Gaussian model we can obtain the entanglement entropy exactly via the covariance matrix that describes the Gaussian ground state of each disordered Hamiltonian. For the same disordered Hamiltonian, we can find an approximation of the ground state via the DMRG, and then calculate the entanglement entropy by manipulating the MPS ground state accordingly.
In Fig.9, we draw the averaged entanglement entropy over the subsystem size for disordered Gaussian Hamiltonians, generated using the couplings of the 3rd inflation step of the $\{3,7\}$-hyperbolic lattice. As we can see, the DMRG method approximates the ground state of the Gaussian models with great accuracy for bond dimension $D=1024$.
Figure 9: We benchmark the DMRG ground state approximation with the exact result out of the covariance matrix formalism, by calculating their averaged entanglement entropy scaling over subsystem size for the Gaussian XX-Heisenberg model ($\Delta_{0}=0$) and disorder cases $r=\{0.4,0.6,0.8,1\}$.
@@ Appendix C Alternative measure of disorder
In Fig. 8(a), we plot $S_{\text{avg}}(\ell=29)$ as a function of the disorder strength for the MQA derived from the inflation rules of the $\{3,7\}$, $\{6,4\}$, $\{4,5\}$, and $\{3,8\}$ hyperbolic lattices (see Table. 1 for the specific inflation rules). For all these case of “holographic” disorder, we can distinguish plateaus of different widths that expand in the disordered regime. We contrast this behavior with non-MQA aperiodic sequences: In Fig. 8(b), we plot $S_{\text{avg}}(\ell=29)$ for the Thue-Morse aperiodic sequence ($a\mapsto ab,b\mapsto ba$) as well as Aubry-André couplings, for both of which it decreases monotonically with increasing $R$. Interestingly, if we draw couplings randomly from a Gaussian distribution (also shown in Fig. 8(b)), the entanglement entropy again decreases monotonically with $R$, without any signs of a disordered critical regime. Specifically, we drew couplings $J_{i}$ from a normal distribution with mean $\mu=10$ and variance $\sigma^{2}\in[0.25,3]$. Each data point corresponds to a fixed $\{\mu,\sigma^{2}\}$ with 2000 random samples, with their average disorder strength $R$ and entanglement $S_{\text{avg}}(\ell=L/2)$ being plotted. These results show that the MQA plateau behavior is very non-generic for random or even quasiperiodic disorder. Of course, any single random sample may deviate sharply from the sample average, in special cases producing high entanglement even for strong disorder.

### v2

@@ Acknowledgments
We wish to thank Lennart Bittel and Tommaso Guaita for fruitful discussions on Gaussian states, Jens Eisert for helpful comments on boundary conditions in critical spin chains, and Zoltán Zimborás for relevant discussions on aperiodic disorder. For all tensor network numerical simulations done in this paper, we used the QSpace tensor library [42, 43, 44] which can exploit general non-Abelian symmetries in tensor network algorithms. The authors are grateful for support from the Einstein Research Unit “Perspectives of a quantum digital transformation”.
@@ Appendix A Covariance matrix formalism and entanglement entropy calculation
A fermionic Gaussian state [26] can be defined as any state
@@
The above expression reveals that the fermionic Gaussian state $\rho$ is completely characterized by the occupation numbers $\langle\tilde{c}^{\dagger}_{k}\tilde{c}_{k}\rangle$ and as a consequence by the correlators $\langle c_{i}^{\dagger}c_{j}\rangle$ and $\langle c_{i}c_{j}\rangle$ after we apply the reverse transformation $c_{k}=U^{\dagger}\tilde{c}_{k}$. It is convenient to study a Gaussian model in the Majorana fermion picture. In this case the Gaussian fermionic state $\rho$ is completely characterized by the covariance matrix
@@
where the density matrix is thermal and $\gamma_{k}$ are the Majorana operators with $\{\gamma_{j},\gamma_{k}\}=2\delta_{j,k}$. The covariance matrix $\Gamma$ is a real, skew-symmetric matrix, which makes its diagonalization practical: For any real, skew-symmetric $2N\times 2N$ matrix $M$, there exists a real special orthogonal matrix $O\in SO(2N)$, such that $M=O\tilde{M}O^{\text{T}}$, where $\tilde{M}$ is a block-diagonal matrix of the form
@@
where $\mu_{j}\in\big{[}-\frac{1}{2},\frac{1}{2}\big{]}$. The eigenvalues $\epsilon_{k}$ of the Hamiltonian $H$ are then given by
@@
Having obtained the eigenvalues $\mu_{j}$, the von Neumann entanglement entropy is given by
$S(\rho)=\sum_{k=1}^{N}\big{[}\nu_{k}\ln(\nu_{k})+(1-\nu_{k})\ln(1-\nu_{k})\big{]}\ ,$ (18)
where $\nu_{k}=\frac{1}{2}-\mu_{k}$. The block-diagonal form of the covariance matrix makes the calculation of the entanglement entropy of subsystems of any length $\ell$ quite convenient.
Figure 8: We benchmark the DMRG ground state approximation with the exact result out of the covariance matrix formalism, by calculating their averaged entanglement entropy scaling over subsystem size for the Gaussian XX-Heisenberg model ($\Delta_{0}=0$) and disorder cases $r\in\{0.4,0.6,0.8,1\}$.
Figure 9: Site-averaged entanglement entropy $S_{\text{avg}}$ for subsystems of size $\ell=29$ over disorder strength for (a): hyperbolic sequences and (b): non-hyperbolic aperiodic sequences. In both (a) and (b) we present the result of the $\{3,7\}$ hyperbolic lattice with a black line for comparison. (c): Averaged entanglement entropy for subsystems of size $\ell=29$ over disorder strength for the Fibonacci and Silver-mean aperiodic sequences. (d): We present the behavior of the averaged entanglement over disorder strength for the 6th inflation step of the $\{3,7\}$ hyperbolic sequence and the 7th inflation step of the Silver-mean sequences. Each represents the asymptotic behavior of the averaged entanglement entropy at a large number of inflation steps for these two cases.
@@ Appendix B Benchmarking DMRG
We can check the accuracy of the DMRG method we used for the disordered non-Gaussian models in our analysis above by applying it also to the Gaussian model. For the Gaussian model we can obtain the entanglement entropy exactly via the covariance matrix that describes the Gaussian ground state of each disordered Hamiltonian. For the same disordered Hamiltonian, we can find an approximation of the ground state via DMRG, and then calculate the entanglement entropy by manipulating the MPS ground state accordingly.
In Fig. 8, we plot the averaged entanglement entropy over the subsystem size for disordered Gaussian Hamiltonians, generated using the couplings of the 3rd inflation step of the $\{3,7\}$ hyperbolic lattice. As we can see, the DMRG method approximates the ground state of the Gaussian models with great accuracy for bond dimension $D=1024$.
@@ Appendix C Alternative measure of disorder
In Fig. 9(a), we plot $S_{\text{avg}}(\ell=29)$ as a function of the disorder strength for the MQA derived from the inflation rules of the $\{3,7\}$, $\{6,4\}$, $\{4,5\}$, and $\{3,8\}$ hyperbolic lattices (see Table 1 for the specific inflation rules). For all these case of “holographic” disorder, we can distinguish plateaus of different widths that expand in the disordered regime. We contrast this behavior with non-MQA aperiodic sequences: In Fig. 9(b), we plot $S_{\text{avg}}(\ell=29)$ for the Thue-Morse aperiodic sequence ($a\mapsto ab,b\mapsto ba$) as well as Aubry-André couplings, for both of which it decreases monotonically with increasing $R$. Interestingly, if we draw couplings randomly from a Gaussian distribution (also shown in Fig. 9(b)), the entanglement entropy again decreases monotonically with $R$, without any signs of a disordered critical regime. Specifically, we drew couplings $J_{i}$ from a normal distribution with mean $\mu=10$ and variance $\sigma^{2}\in[0.25,3]$. Each data point corresponds to a fixed $\{\mu,\sigma^{2}\}$ with 2000 random samples, with their average disorder strength $R$ and entanglement $S_{\text{avg}}(\ell=L/2)$ being plotted. These results show that the MQA plateau behavior is very non-generic for random or even quasiperiodic disorder. Of course, any single random sample may deviate sharply from the sample average, in special cases producing high entanglement even for strong disorder.